OpenStudy (anonymous):

a1=0.11,a5=895.956875,r=9.5 help please

OpenStudy (anonymous):

Geometric sequence? an = a1*r^(n-1) Given a1 = 0.11 and r=9.5 you can construct the general term using this formula: an = 0.11*(9.5)^(n-1) so a5 = 0.11*(9.5)^4 = 895.956875